OpenStudy (anonymous):

2x2 – 4x = 0 A. {0, 2} B. {–6, –2} C. {0, –2} D. {0, }

OpenStudy (anonymous):

D. (0,1/2)

jagr2713 (jagr2713):

First we factor out the common term 2x an get 2x(x-2)=0 When will x(x−2) equal zero?

OpenStudy (anonymous):

okay! @jagr2713

jagr2713 (jagr2713):

When will x(x−2) equal zero? x=0 and???

OpenStudy (anonymous):

im confused?

jagr2713 (jagr2713):

ok we factored the commen term right? wich is 2x so 4x/2 is 2 2x(x-2)=0 now x(x-2) equals zeo when x=0 or x-2=0. We get the x-2 in the equation ok now we have one part of the answer which is (0,?) solve for x-2=0

jagr2713 (jagr2713):

zero*

OpenStudy (anonymous):

2

jagr2713 (jagr2713):

Good so the answer is?

OpenStudy (anonymous):

A?

jagr2713 (jagr2713):

Correct :D